# Visual place recognition for aerial imagery: A survey

### v1

@@ 1 Introduction
To date, reviews and comparisons of VPR systems have primarily focused on indoor and outdoor data, neglecting satellite and aerial imagery (40; 22; 59; 8). This paper aims to deliver a thorough comparative analysis of current VPR and re-ranking methodologies in the context of aerial imagery, as well as to identify the most effective map representation for this application. To achieve this, we introduce a novel methodology for constructing databases for the VPR task, which encompasses the determination of the optimal scale level and the level of overlap between adjacent map tiles. Additionally, we present an open-source benchmark that incorporates a range of popular VPR and re-ranking techniques.
@@ 2 Related Work
This section provides a comprehensive review of existing Visual Place Recognition systems, highlighting various re-ranking techniques designed to enhance their efficacy (4). It particularly emphasizes on the review articles pertinent to this topic. Special consideration is given to the VPR in aerial imagery, detailing commonly utilized datasets. Additionally, this section discusses the overview of evaluation technologies employed in assessing the performance of VPR methodologies.
@@ 2.1 Visual Place Recognition methods
The goal of VPR algorithms is to find the most appropriate image among a database of images and the locations that match them. This objective constitutes a subset within the broader domain of Image Retrieval (17; 55), with applications spanning search engines, visual content analysis, and robotics. The classic approach for Visual Place Recognition, as employed in early works (14; 21), involves utilizing descriptors, which are compact representations of images. They are obtained by aggregating local features of images into a vector of fixed length. Then, L2 or cosine similarity is used to search in the database of descriptors.
During the nascent phases of feature aggregation methodology, the bag-of-words (21) technique gained significant traction. Subsequent to this, the scientific community was acquainted with the VLAD (Vector of Locally Aggregated Descriptors) (3) methodology, which entails the fusion of local descriptors obtained from keypoints within an image into a vectorized form via the calculation of disparities between each descriptor and a set of predefined visual clusters. Following this, a trainable adaptation of VLAD, denoted as NetVLAD (2), was introduced to the community. NetVLAD utilizes AlexNet (31) and VGG-16 (51) as base architectures. More contemporary VPR techniques like CosPlace (6) and its viewpoint-robust variant, EigenPlaces (9), are evaluated on both VGG-16 and the newer backbone ResNet (26), with the latter demonstrating superior performance. MixVPR (1) adopts ResNet as its backbone and employs the all-MLP feature aggregation technique.
@@ 2.2 Visual Place Recognition surveys
Over the past few years, a number of surveys exploring different approaches to Visual Place Recognition have become available to the community. The purpose of some (5; 40) is only to review different approaches to VPR or to find possible applications (22), and no data-driven comparisons are made.
@@ 2.3 Re-ranking methods
Another approach for the re-ranking task is to use dense local features. For example, Patch-NetVLAD (25) applies VLAD (3) aggregation to extract image patch descriptors and match them using RANSAC or Rapid Scoring. Recent SOTA solutions in this area are R2Former (63) and SelaVPR (37). These methods use ViT backbones to extract dense local features both for their matching directly and for further aggregation. Thus, these solutions combine both classical Image Retrieval approach and re-ranking.
@@ 3 Methodology
The suggested approach adheres to a typical framework for Visual Place Recognition, with its comprehensive workflow depicted in Figure 2. The process is divided into two main segments: offline and online. The offline segment involves building a database for VPR techniques. The online segment includes the Image Retrieval component as well as Local Alignment, which determines the precise position of the system using the chosen frame. Although our approach shares similarities with other VPR benchmarks, it incorporates several unique innovations. Although the challenge of determining appropriate zoom and overlap levels between frames in database construction from aerial imagery has been previously highlighted in the academic community (16; 24), we introduce a generalized parameterization method to address this problem. Additionally, we propose two new metrics that are better suited for aerial imagery compared to the widely used Recall@k metric.
@@ 3.2 Image retrieval
Following the standard scheme for the Visual Place Recognition task, we use pre-calculated descriptors of map tiles, as well as local features that are needed for the re-ranking step. The process of query frame localization is as follows. For a frame, its descriptor is computed using VPR method. Then the database is searched from the descriptors and $N$ nearest frames are selected. Using $N$ candidate frames, $K$ best matches are selected using re-ranking algorithms.
@@ 3.3 Local alignment
The task of Local Alignment is to calculate the geo-coordinates of the query image using the previously found corresponding satellite map tile. To do this, we use the keypoints of the query and satellite images and match them. Then we compute the perspective transformation, which allows us to map the query image to the satellite map tile. Since the satellite image itself is a rectangle with previously known geo-coordinates of its corners, it is possible to convert pixel coordinates to geographic coordinates. For this purpose, our approach provides Web Mercator model, which is used in Google Maps.
@@ 3.5 Metrics
In the visual place recognition task, the standard metric is Recall@k (8; 2; 7; 10; 25; 28; 35; 43; 44; 56), which measures the percentage of query images for which at least one of the $K$ best images selected by the method is no further away than a given threshold. This metric assumes that the geo-coordinates of the images retrieved from the database are utilized as the resultant geo-coordinates. However, within the realm of aerial data, this approach isn’t appropriate due to the substantial size of map tiles, raising ambiguity regarding the association of specific geo-coordinates with images from the database. To address this challenge, two adaptations of the original Recall@k metric have been suggested. One assesses the complete visual geolocalization process, encompassing the local alignment stage, while the other focuses solely on the global localization phase. It’s worth noting that for both suggested metrics, we assume that the query images are nadir images, and we consider their 2D projection. Consequently, we presume that the coordinates of the query image represent the latitude and longitude of its center.
@@ 4.4 Discussion
In terms of temporal and memory metrics, a clearer pattern emerges. SelaVPR (37) emerges as the swiftest method across most stages, although certain other methodologies demonstrate comparable speeds at select stages. Additionally, SelaVPR demands the least memory allocation for storing global descriptors among all methods. However, SelaVPR necessitates a higher memory allocation for storing local features compared to SuperPoint (18), the feature descriptor employed by SuperGlue (48) and LightGlue (34).
@@ 5 Conclusion
This paper presents a comprehensive evaluation of different VPR and re-ranking methodologies tailored to aerial imagery. By introducing a novel approach for database construction, we provide a methodology for evaluating VPR methods in the context of aerial data. Furthermore, our open-source benchmark facilitates the combination and comparison of various VPR and re-ranking techniques. This paper serves as a starting point for researchers and practitioners exploring the application of aerial VPR algorithms, emphasizing the significance of hyperparameter optimization (specifically zoom and overlap) in the generation of map tiles. Additionally, it engages in discourse concerning various facets of state-of-the-art localization algorithms within the aerial domain. We have made all contributions, including code, public.

### v2

@@ 1 Introduction
To date, reviews and comparisons of VPR systems have primarily focused on indoor and outdoor data, neglecting satellite and aerial imagery (43; 24; 63; 9). This paper aims to deliver a thorough comparative analysis of current VPR and re-ranking methodologies in the context of aerial imagery, as well as to identify the most effective map representation for this application. To achieve this, we introduce a novel methodology for constructing databases for the VPR task, which encompasses the determination of the optimal scale level and the level of overlap between adjacent map tiles. Additionally, we present an open-source benchmark that incorporates a range of popular VPR and re-ranking techniques.
@@
The paper is structured to address the following components: (Section 2) reviews related work on VPR methods, re-ranking techniques, and datasets for aerial imagery. (Section 3) describes the methodology, including database construction and evaluation metrics. (Section 4) and (Section 5) cover the experimental setup, results, and discussion.
@@ 2 Related Work
This section provides a comprehensive review of existing Visual Place Recognition systems, highlighting various re-ranking techniques designed to enhance their efficacy (5). The summary of this review is provided in Table 1. It particularly emphasizes on the review articles pertinent to this topic. Special consideration is given to the VPR in aerial imagery, detailing commonly utilized datasets. Additionally, this section discusses the overview of evaluation technologies employed in assessing the performance of VPR methodologies.
Ref Method Year Backbone Test domains
(16) FAB-MAP 2008 Urban
(23) Gálvez-López et al. 2012 Indoor, Urban
(3) NetVLAD 2016 AlexNet, VGG-16 Urban
(53) Shetty et al. 2019 AlexNet Aerial
(42) Mantelli et al. 2019 Aerial
(64) CoHOG 2020 Indoor, Urban
(71) Zhuang et al. 2021 ResNet Aerial
(17) Dai et al. 2021 ViT-S Aerial
(12) Bianchi et al. 2021 Custom autoencoder Aerial
(7) CosPlace 2022 VGG-16, ResNet Urban
(10) Eigenplaces 2023 VGG-16, ResNet Urban
(1) MixVPR 2023 ResNet Urban
(31) AnyLoc 2023 DINOv2 Indoor, Urban, Underwater,
Aerial, Subterranean
(29) SALAD 2023 DINOv2 Urban
(59) Wang et al. 2024 ResNet Aerial
(40) SelaVPR 2024 DINOv2 Urban
Table 1: Overview of existing Visual Place Recognition systems
@@ 2.1 Visual Place Recognition methods
The goal of VPR algorithms is to find the most appropriate image among a database of images and the locations that match them. This objective constitutes a subset within the broader domain of Image Retrieval (19; 58), with applications spanning search engines, visual content analysis, and robotics. The classic approach for Visual Place Recognition, as employed in early works (16; 23), involves utilizing descriptors, which are compact representations of images. They are obtained by aggregating local features of images into a vector of fixed length. Then, L2 or cosine similarity is used to search in the database of descriptors.
During the nascent phases of feature aggregation methodology, the bag-of-words (23) technique gained significant traction. Subsequent to this, the scientific community was acquainted with the VLAD (Vector of Locally Aggregated Descriptors) (4) methodology, which entails the fusion of local descriptors obtained from keypoints within an image into a vectorized form via the calculation of disparities between each descriptor and a set of predefined visual clusters. Following this, a trainable adaptation of VLAD, denoted as NetVLAD (3), was introduced to the community. NetVLAD utilizes AlexNet (33) and VGG-16 (54) as base architectures. However, methods using hand-crafted features have also seen development. For example, Zaffar et al. (64) employ Histogram-of-Oriented-Gradients (HOG) descriptors and image entropy to extract regions of interest and perform regional-convolutional descriptor matching. More contemporary VPR techniques like CosPlace (7) and its viewpoint-robust variant, EigenPlaces (10), are evaluated on both VGG-16 and the newer backbone ResNet (28), with the latter demonstrating superior performance. MixVPR (1) adopts ResNet as its backbone and employs the all-MLP feature aggregation technique.
@@
One of the principal drawbacks of numerous works in the field of Visual Place Recognition is the insufficient attention given to the performance issues of the proposed methods. Concurrently, there is a trend towards utilizing large foundation models, such as DINOv2. For instance, AnyLoc employs the largest DINOv2 model in terms of parameters—ViT-G. Although this model enhances prediction quality, the time required for prediction can be excessively long, even on high-performance devices. Therefore, it is anticipated that future research in VPR will increasingly focus on addressing performance issues, particularly in the context of embedded devices.
@@ 2.2 Visual Place Recognition surveys
Over the past few years, a number of surveys exploring different approaches to Visual Place Recognition have become available to the community. The purpose of some (6; 43) is only to review different approaches to VPR or to find possible applications (24), and no data-driven comparisons are made.
@@
Li et al. (35) evaluate fifteen global descriptor methods for VPR by analyzing their performance across six diverse datasets, encompassing indoor environments, urban roads, suburban areas, and natural scenery. The study provides design recommendations for enhancing global descriptors and examines the trade-offs between matching performance and computational efficiency in practical VPR applications. However, this study does not address the application of VPR techniques in the context of aerial imagery.
@@ 2.3 Re-ranking methods
Another approach for the re-ranking task is to use dense local features. For example, Patch-NetVLAD (27) applies VLAD (4) aggregation to extract image patch descriptors and match them using RANSAC or Rapid Scoring. Recent SOTA solutions in this area are ETR (66), R2Former (69) and SelaVPR (40). These methods use ViT backbones to extract dense local features both for their matching directly and for further aggregation. Thus, these solutions combine both classical Image Retrieval approach and re-ranking.
@@ 2.4 VPR in aerial images
Wang et al. (59) identify weather and lighting variations as significant challenges for visual geo-localization systems. To mitigate these issues, they propose MuSe-Net, an adaptive learning framework incorporating a dual-path CNN to minimize environmental style discrepancies and a Residual SPADE module to enhance feature discrimination and optimize training.
@@ 3 Methodology
The suggested approach adheres to a typical framework for Visual Place Recognition, with its comprehensive workflow depicted in Figure 2. The process is divided into two main segments: offline and online. The offline segment involves building a database for VPR techniques. The online segment includes the Image Retrieval component as well as Local Alignment, which determines the precise position of the system using the chosen frame. Although our approach shares similarities with other VPR benchmarks, it incorporates several unique innovations. Although the challenge of determining appropriate zoom and overlap levels between frames in database construction from aerial imagery has been previously highlighted in the academic community (18; 26), we introduce a generalized parameterization method to address this problem. Additionally, we propose two new metrics that are better suited for aerial imagery compared to the widely used Recall@k metric.
@@ 3.1 Database construction
In our proposed framework, the capability to generate maps with arbitrary overlap and zoom levels is implemented using the OpenCV (13) library. The tool requires only a pre-downloaded map with zero overlap and any specified zoom level. A script for downloading maps using the Google Maps API is also available within our framework.
@@ 3.2 Image retrieval
Following the standard procedure for the Visual Place Recognition (VPR) task, we utilize pre-computed descriptors of map tiles and local features required for the re-ranking step. The query frame localization process is as follows: for each frame, its descriptor is computed using the VPR method. In our study, we employed the authors’ implementations based on PyTorch (2). The database is then searched using these descriptors, and the $N$ nearest frames are identified. Our tool leverages the Faiss (21) library for efficient descriptor database search. Subsequently, among the $N$ candidate frames, $K$ best matches are selected using re-ranking algorithms.
@@ 3.3 Local alignment
The task of Local Alignment is to calculate the geo-coordinates of the query image using the previously found corresponding satellite map tile. To do this, we use the keypoints of the query and satellite images and match them. Then we compute the perspective transformation, which allows us to map the query image to the satellite map tile. Since the satellite image itself is a rectangle with previously known geo-coordinates of its corners, it is possible to convert pixel coordinates to geographic coordinates. For this purpose, our approach provides Web Mercator model, which is used in Google Maps. Functionality for local alignment is implemented in the proposed framework using the corresponding methods cv2.findHomography and cv2.perspectiveTransform from the OpenCV library (13).
@@ 3.5 Metrics
In the visual place recognition task, the standard metric is Recall@k (9; 3; 8; 11; 27; 30; 38; 46; 47; 60), which measures the percentage of query images for which at least one of the $K$ best images selected by the method is no further away than a given threshold. This metric assumes that the geo-coordinates of the images retrieved from the database are utilized as the resultant geo-coordinates. However, within the realm of aerial data, this approach isn’t appropriate due to the substantial size of map tiles, raising ambiguity regarding the association of specific geo-coordinates with images from the database. To address this challenge, two adaptations of the original Recall@k metric have been suggested. One assesses the complete visual geolocalization process, encompassing the local alignment stage, while the other focuses solely on the global localization phase. It’s worth noting that for both suggested metrics, we assume that the query images are nadir images, and we consider their 2D projection. Consequently, we presume that the coordinates of the query image represent the latitude and longitude of its center.
@@ 4.4 Discussion
In terms of temporal and memory metrics, a clearer pattern emerges. SelaVPR (40) emerges as the swiftest method across most stages, although certain other methodologies demonstrate comparable speeds at select stages. Additionally, SelaVPR demands the least memory allocation for storing global descriptors among all methods. However, SelaVPR necessitates a higher memory allocation for storing local features compared to SuperPoint (20), the feature descriptor employed by SuperGlue (51) and LightGlue (37). When addressing time and memory constraints in the scaling of VPR methods, the principal challenge is the substantial memory consumption associated with local features. For large-scale maps, a feasible solution is to use global localization methods without re-ranking step. Among the VPR methods we assessed, all are potentially suitable for this approach, with the exception of AnyLoc, which is excluded due to the high dimensionality of its descriptor.
It is important to acknowledge that while a few visual geolocalization techniques demonstrate frame rates of 1 FPS or higher, which may be adequate for various practical applications, our experiments were conducted using a high-performance computing system. Consequently, evaluating the performance of these methods on microcomputers integrated into airborne devices represents a logical extension of our research.
@@ 5 Conclusion
This paper presents a comprehensive evaluation of different VPR and re-ranking methodologies tailored to aerial imagery. By introducing a novel approach for database construction, we provide a methodology for evaluating VPR methods in the context of aerial data. Furthermore, our open-source benchmark facilitates the combination and comparison of various VPR and re-ranking techniques. This paper serves as a starting point for researchers and practitioners exploring the application of aerial VPR algorithms, emphasizing the significance of hyperparameter optimization (specifically zoom and overlap) in the generation of map tiles. Additionally, it engages in discourse concerning various facets of state-of-the-art localization algorithms within the aerial domain. Moreover, the results of our study have potential applications in the deployment of visual geolocalization systems on real-world airborne platforms. While our methodology alone may not provide comprehensive robustness, it can be effectively augmented with additional sensors, such as inertial measurement units (IMUs). This integration enhances its utility for Visual Inertial Odometry (VIO) and Simultaneous Localization and Mapping (SLAM) systems, particularly for periodic location refinement and loop closure tasks. Additionally, our methodology could serve as a dependable emergency localization fallback in the event of an unexpected GNSS signal loss. We have made all contributions, including code, public.